# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

### v1

@@ VII-A Privacy and Utility Trade-off
One of the most critical challenges in network traffic synthesis is ensuring privacy while maintaining the utility of the generated data. Real network traffic often contains sensitive information such as IP addresses, communication patterns, and user behavior, making privacy protection essential. However, strong privacy preservation measures, such as differential privacy, often degrade the quality of the synthetic data, leading to reduced performance in downstream tasks. This opens up to the privacy-utility trade-off.
@@
Differential Privacy introduces noise into the data to mask individual entries, but this can distort key traffic patterns. The challenge is to find the right balance between the level of noise (controlled by the privacy parameter $\epsilon$) and the ability of the synthetic data to retain useful properties for tasks like anomaly detection or traffic analysis. Fine-tuning $\epsilon$ requires a deep understanding of the privacy risks and application requirements. They can lead to a loss of granularity in the data, making it difficult to maintain fine-grained traffic features that are essential for some applications. Some preliminary work has been done in [171, 87] to address the privacy utility trade-off by incorporating DP into the GAN generation. Authors of [104, 28] have tried to address the membership inference attacks on GANs.
VII-B Adhering to Protocol Standards
@@
Developing synthetic network data that satisfies timestamp and protocol requirements is essential, as networks operate under strict protocol standards that govern data integrity and transmission quality. Ensuring that synthetic traffic adheres to these protocols is critical for realistic and useful synthetic data. Over 80% of the work does not consider developing network traffic generation that adheres to protocol standards. This has been addressed to a certain level in [64, 65] by integrating ControlNet into DM for data generation to adhere to protocol standards and performing a post-processing step to correct the images to ensure that they adhere to the protocol. However, there still exist certain challenges to be addressed.
VII-B1 Protocol Compliance in Data Generation
@@
Network protocols define limits on packet sizes, acceptable payload structures, header formats, and timing between packets. For instance, protocols like TCP/IP and UDP have maximum packet size constraints and specific sequencing requirements. Synthetic data must comply with these standards to ensure usability for applications like simulation and stress testing of network systems.
VII-B2 Timestamp Synchronization and Temporal Coherence
@@
Network data is highly time-sensitive, with applications that rely on precise timestamps and sequence orders for accurate simulations. Synthesis algorithms must handle timestamps and ensure that events in the synthetic dataset mirror the temporal coherence found in real network traffic. This challenge becomes even more pronounced in applications like intrusion detection, where real-time analysis of packet sequences is essential.
@@ VII-C Capturing Complex Network Behaviors
Real-world network traffic exhibits highly complex behaviors, shaped by dynamic routing, load balancing, and bursty traffic flows. Capturing these nuances in synthetic data remains a significant challenge. Some efforts have addressed this to an extent. For instance, the authors of [68, 94, 137] employ packet binning to condense time series traces while preserving essential feature spaces for long traces. Similarly, Sivaroopan et al. [137, 136] leverage a GASF imaging mechanism to capture temporal correlations in packet-level features. However, despite these advancements, further challenges remain in accurately modeling the intricate characteristics of network traffic.
VII-C1 Traffic Diversity
Network traffic varies greatly depending on the type of applications, devices, and protocols in use. Synthetic data generators need to be flexible enough to capture this diversity, from low-latency traffic such as video streaming to high-throughput traffic such as large data transfers. However, existing synthesis methods often fail to replicate these variations effectively, leading to over-simplified or unrealistic synthetic traffic patterns.
VII-C2 Temporal Dynamics and Seasonality
Network traffic is not static; it changes over time due to user behavior and network conditions. Ensuring that synthetic data models these temporal dynamics, including short-term fluctuations and long-term seasonality, is a major challenge.
VII-C3 High-dimensional Feature Spaces
Network traffic data is often high-dimensional, encompassing a wide range of features such as packet sizes, inter-arrival times, protocol flags, and source-destination IP pairs. Capturing correlations and dependencies across these features in a synthetic dataset is difficult. Synthetic data may either over-simplify these relationships, lose important interactions, or become too complex to be useful for downstream tasks.
@@ VII-D Evaluating Synthetic Data Quality
Another key challenge lies in developing standardized metrics for evaluating the quality of synthetic network traffic data. While we have discussed several metrics such as statistical similarity and performance on downstream tasks, there is no one-size-fits-all solution. There are many papers  [69, 67, 137, 136] attempting to provide high-quality synthetic data, however, there are challenges yet to be solved.
VII-D1 Lack of Benchmark Datasets
Unlike other fields such as computer vision or natural language processing, there are relatively few publicly available, standardized benchmark datasets for synthetic network traffic evaluation. This makes it difficult to compare the performance of different synthesis approaches, especially for privacy-preserving techniques.
VII-D2 Generalization vs. Overfitting
Striking the right balance between generalization and overfitting in synthetic data generation remains a major issue. Overfitting to the real traffic can lead to privacy risks, while excessive generalization may result in synthetic data that fails to capture the essential characteristics of the network traffic, reducing its utility.
VII-D3 Task-Specific Evaluation
The effectiveness of synthetic data can vary significantly depending on the specific task it is used for, such as traffic classification, anomaly detection, or intrusion prevention. This calls for task-specific evaluation metrics, but defining these metrics can be complex and may require different evaluation frameworks for each application.
@@ VII-E Computational Complexity and Scalability
As networks expand in size and complexity, the computational demands of generating and evaluating synthetic data grow accordingly. This also poses a particular challenge for real-time network traffic generation on edge devices, where computational resources are limited. To address this, the authors of [83] introduced lightweight DMs designed for efficient data generation on edge devices. However, further challenges remain in balancing computational efficiency with the accuracy and fidelity of the generated traffic.
VII-E1 Scalability of Data Generation
@@
Large-scale networks generate massive amounts of traffic data, and synthesizing data at this scale requires significant computational resources. Many existing synthesis methods struggle to scale effectively, leading to bottlenecks in generating synthetic data for large network environments.
VII-E2 Real-time Synthesis
@@
In some applications, synthetic data may need to be generated in real-time to support testing and analysis under current network conditions. Achieving real-time synthesis is particularly challenging when attempting to replicate complex traffic patterns while maintaining high fidelity and privacy guarantees.
In summary, network traffic synthesis has tremendous potential but faces significant challenges. These challenges include preserving privacy without compromising data utility, capturing complex network behavior accurately, and developing robust evaluation metrics. Additionally, scalability remains an issue, as generating and evaluating synthetic data for large networks requires extensive resources. Ensuring that synthetic data aligns with protocol standards and accurately represents both typical and unseen network scenarios will further enhance the realism and applicability of synthetic traffic data for testing, simulation, and security analysis. Overcoming these hurdles is essential for the broad adoption of synthetic network traffic data across varied applications.
VIII Future Directions
As communication networks evolve to meet increasing demand and diverse operational environments, accurately modeling and replicating real-world network conditions will ensure robust performance, efficient resource allocation, and seamless user experiences. In this context, network traffic generation is expected to become crucial in optimizing network services.
One such practical application of network traffic generation can be illustrated through Ericsson’s 5G Digital Twin [110], simulated in NVIDIA Omniverse Enterprise. Ericsson employs Digital Twins (DTs) i.e virtual replicas of physical systems, to model the interaction between 5G cells and urban environments in real time. While these models can predict coverage based on the physical environment, they currently lack the ability to simulate real network conditions based on factors such as variations in traffic volume, application-specific patterns, and congestion dynamics. This is where network traffic generation becomes transformative. Creating synthetic traffic that accurately mimics real-world network behavior, enables researchers and engineers to simulate, test, and refine network performance under realistic conditions. Integrating traffic-aware simulations into 5G/6G Digital Twins will significantly enhance predictive accuracy, optimize network configurations, and improve overall service quality. Beyond twins, one may also envision the possibility of networks of synthetic traffic sources ( a digital twin without a physical counterpart) generating and inserting traffic to physical networks for tasks such as camouflaging or protecting privacy.
In the remainder of this section, we explore potential enhancements and future directions for improving network traffic generation.
VIII-A Commercialization and Multimodality of Emerging Deep Generative Models
Future commercial network traffic generators could significantly benefit from the integration of advanced generative techniques such as DMs and Transformer-based architectures (GPTs), which have set new standards in generative modeling. GPTs, being pre-trained on extensive datasets as large language models, exhibit unparalleled sequence generation capabilities. This knowledge aligns seamlessly with the requirements of time-series generation paradigms, such as network traffic data, where temporal and sequential dependencies are critical. By bridging the gap between the LLM’s word embedding space and the embedding space of time-series tokens or packet flows, GPTs can be fine-tuned to generate realistic and contextually rich network traffic patterns. Similarly, DMs, known for their iterative refinement of data from noise, offer an exceptional ability to model intricate data distributions and generate synthetic data with high fidelity. These models could be particularly effective in mimicking the statistical and temporal properties of network traffic, capturing both short-term variations and long-term dependencies.
Beyond these advancements, the future of network traffic generation lies in multimodal deep generative models that integrate multiple data types, including textual descriptions, images, structured network metadata (e.g., packet headers, flow statistics), and raw RF signals. DMs such as Stable Diffusion can be conditioned on network topology diagrams or textual descriptions of attack scenarios to generate structured protocol-compliant synthetic traffic. Similarly, transformer (e.g. Contrastive Language-Image Pretraining (CLIP) [125]) models can align textual representation of network behavior with images or other modalities, allowing for more context-aware traffic synthesis that is more interpretable and adaptive. GPTs can extend their sequence modeling capabilities beyond packet flows by incorporating visual and textual network telemetry, enabling cross-modal learning for real-time network traffic simulation. The synergy of these techniques in commercial applications could enable scalable and precise traffic generation solutions, addressing long-standing challenges such as data imbalance, privacy preservation, and real-time traffic simulation. By leveraging these SOTA multimodal generative models, the next generation of commercial traffic generators could offer enhanced capabilities, transforming how synthetic network traffic is created and utilized for cybersecurity, network optimization, and anomaly detection.
VIII-B Synthetic Data Generation for Unseen Network Scenarios
As discussed earlier, synthesizing network data that represents unseen or hypothetical scenarios is crucial for evaluating system resilience and preparedness in various environments.
@@ VIII-B Synthetic Data Generation for Unseen Network Scenarios
Generative models can simulate unseen network conditions by learning from available datasets and extrapolating patterns for unobserved scenarios, such as extreme network loads, rare cyberattacks, or unusual traffic distributions in geographically isolated locations. Additionally, when real-world data for specific platforms or protocols is unavailable, synthetic data generation can incorporate domain knowledge to approximate these environments. However, the key challenge lies in ensuring that the generated data remains both valid and useful for downstream tasks.
One possible strategy is to follow a similar approach as Neural Style Transfer (NST) [51] which combines the content of one image with the artistic style of another. NST is based on the intuition that the content (structure and high-level features) and style (textures, patterns, and colors) of an image can be extracted using different layers of a deep convolutional network and that optimizing a new image to match the content features of one image and the style features of another generates a result that retains the structure of the content image while adopting the artistic characteristics of the style image. To follow a similar approach, we first define two attributes for each network trace, Content which refers to the actual content/payload carried by the traffic trace (such as a video stream, or website), and Domain which refers to the conditions of the network which affects the traffic. For instance, when considering a network trace for a user streaming videos, the specific video title would be the content and the domain could be the specific network the user is connected to (wired or wireless) or the specific geographical location. The notation $T_{n}(C_{x},D_{y})$ represents a traffic trace corresponding to content $x$ and domain $y$. For example, consider a scenario where traffic traces with content $C_{1}$ from domain $D_{1}$, and traffic traces with content $C_{2}$ from domain $D_{2}$ are available, but traffic traces with content $C_{2}$ from domain $D_{1}$ is not available and needs to be generated. Analogous to NST, if we can extract content-specific features from $T_{2}(C_{2},D_{2})$ and domain-specific features from $T_{1}(C_{1},D_{1})$, these features can be combined to generate a new sample $T_{3}(C_{2},D_{1})$ representing $C_{2}$ and $D_{1}$.
Unlike with images where the exact layers of a deep convolutional network to extract the content and style from images are known, to the best of our knowledge, no prior research has addressed a similar approach for network traffic traces, making this an open research problem. However, P. Smolovic et al[138] introduced the use of triplet loss to extract video-specific (content) features from network traffic traces, independent of the platform on which the video was streamed (domain). Building on this, we propose that using a similar methodology leveraging contrastive loss functions, such as triplet loss, to extract both content-specific and domain-specific features from network traffic traces is feasible. Once the content-specific and domain-specific features are extracted, the same approach as NST can be followed to generate the new sample by optimizing a random input to match the extracted content and domain.
An alternative approach would be to leverage neural networks with an encoder-decoder architecture, such as a VAE. These networks can be trained to ensure that the latent space forms distinct clusters corresponding to each domain and content type. For example, as illustrated in Figure 10, this approach can be used to create distinct clusters for each domain, and within each domain, further distinct clusters are formed based on content. Once trained, the latent vector for a traffic trace could be extrapolated in the latent space to generate a trace with a different content or domain. For example, Figure 10 shows a scenario where the model is trained on traces with content $C_{1}$ from both domains and $C_{2}$ from $D_{2}$ only. Accordingly, it will learn both $C_{1}$ and $C_{2}$ clusters in $D_{2}$ and only $C_{1}$ cluster in $D_{1}$. Once trained a trace $T(C_{1},D_{1})$ can be extrapolated in the latent space to create $C_{2}$ in $D_{1}$ to match context $C_{2}$ based on the relationship between $C_{1}$ and $C_{2}$ in $D_{2}$. Then the decoder can generate a realistic network traffic sample $T(C_{2},D_{1})$ that represents context $C_{2}$ and domain $D_{1}$.
@@
IX Conclusion
This survey highlights the importance of synthetic network traffic generation in overcoming challenges associated with real-world data collection, such as privacy concerns, scalability issues, and the need for high-quality data. We reviewed and categorized existing approaches into statistical methods and deep learning-based models, while also examining their extensions into network traffic simulators and commercially available tools. This structured classification provides insights into the strengths, limitations, and practical applications of each approach, illustrating how the field has evolved and how these methods are being used to synthesize different scenarios.
Despite progress, there are still challenges to be addressed. Preserving user privacy, ensuring realistic data that accurately reflect complex protocol behaviors, and improving the usability of synthetic data for tasks like network security and performance optimization remain significant hurdles. Advanced AI models like transformers and diffusion models offer new possibilities for synthesizing unseen networking scenarios.
By offering a detailed taxonomy of data types, network traffic generation methods and unresolved issues, this paper aims to inspire innovative solutions and guide researchers and practitioners in advancing synthetic network traffic generation methodologies for robust and efficient network traffic analysis. As synthetic data generation continues to advance, it holds the potential to transform network traffic analysis, making it more efficient, secure, and adaptable to the ever-growing demands of modern networks.

### v2

@@ VII-A Privacy and Utility Trade-off
Challenge: Tradeoff between privacy and utility is a main challenge in network traffic generation. Real network data often contains sensitive information such as IP addresses, and communication patterns. making privacy protection essential. However, applying mechanisms such as DP, often degrade the quality of the synthetic data, leading to reduced performance in downstream tasks. This opens up the privacy-utility trade-off. DP introduces noise into the data to mask individual entries, but this can distort key traffic patterns. The challenge is to find the right balance between the level of noise (controlled by the privacy parameter $\epsilon$) and the ability of the retaining useful properties in synthetic data. Fine-tuning $\epsilon$ requires a deep understanding of the privacy risks and application requirements. Some preliminary work has been done in [197, 102] to address the privacy-utility trade-off by incorporating DP into the GAN generation, nevertheless requires further improvements.
Future Direction: Future privacy-preserving traffic generators can leverage advanced generative models such as DMs and Transformers, incorporating noise at more context-aware levels (e.g., feature-specific perturbations). Multimodal generative models may allow selective obfuscation of sensitive features while preserving utility in others. Beyond DTs, one may also envision the possibility of networks of synthetic traffic sources ( a DT without a physical counterpart) generating and inserting traffic to physical networks for tasks such as camouflaging or protecting privacy. With the rise of Large Language Models (LLMs) based on GPT architectures across various domains, users increasingly rely on cloud-based models, necessitating the filtering of private data before transmission to ensure privacy and security.
VII-B Adhering to Protocol Standards and Realistic Conditions
@@
Challenge: Synthetic network data has to be representative of the timestamp and volume distributions of underlying networks, protocols and conditions they claim to emulate, and must not be in violation of conditions imposed by the strict protocol standards that govern data integrity and transmission quality. Ensuring that synthetic traffic adheres to the protocols is critical for realistic and useful synthetic data. Over 80% of the work does not consider developing network traffic generation that adheres to protocol standards. This has been addressed to a certain level in [76, 77] by integrating ControlNet into DM for data generation to adhere to protocol standards and performing a post-processing step to correct the images to ensure that they adhere to the protocol. However, there still exist certain challenges to be addressed.
@@
Protocol Compliance: Network protocols define limits on packet sizes, acceptable payload structures, header formats, and timing between packets. For instance, protocols like TCP/IP and UDP have maximum packet size constraints and specific sequencing requirements. Synthetic data must comply with these standards to ensure usability for applications like simulation and stress testing of network systems.
@@
Timestamp Synchronization and Temporal Coherence: Network data is highly time-sensitive, with applications that rely on precise timestamps and sequence orders for accurate simulations. Synthesis algorithms must handle timestamps and ensure that events in the synthetic dataset mirror the temporal coherence found in real network traffic. This challenge becomes even more pronounced in applications like IDS, where real-time analysis of packet sequences is essential.
Further, it is nearly impossible to collect data for different realistic conditions to statisfy every possible scenarios. Network traffic synthesis will be a major player to alleviate this problem. This is further discussed in section VII-F.
Future Direction: Future models could explicitly incorporate protocol rules as conditional constraints or even embed Request for Comments (RFC)-level formal specifications. Additionally, timestamp-aware generative models trained using temporal coherence losses can better preserve the temporal order of packets. These improvements are crucial for generating data that conforms to real-world constraints beyond mere protocol adherence. Moreover, realistic simulation environments such as 5G DTs can serve as testbeds for validating protocol- and time-aware data generation under operational conditions. Embedding traffic generators in AI-based network design frameworks like Ericsson’s 5G DT [127] opens new avenues for training, evaluation, and deployment of realistic synthetic traffic under evolving network dynamics.
@@ VII-C Capturing Complex Network Behaviors
Challenge: Network traffic exhibits highly complex behaviors, shaped by dynamic routing, load balancing, and bursty traffic flows. Capturing these nuances in synthetic data remains a significant challenge. Some efforts have addressed this to an extent.For example, [81, 110, 157] use packet binning to compress time series traces while retaining key features, and [157, 156] time-series to 2D imaging to capture complex temporal correlations in packet-level data. However, despite these advancements, further challenges remain in accurately modeling the intricate characteristics of network traffic.
Traffic Diversity: Network traffic varies greatly depending on the type of applications, devices, and protocols in use. Synthetic data generators need to be flexible enough to capture this diversity, from low-latency traffic such as video streaming to high-throughput traffic such as large data transfers. However, existing synthesis methods often fail to replicate these variations effectively, leading to over-simplified or unrealistic synthetic traffic patterns.
Temporal Dynamics and Seasonality: Network traffic is not static; it changes over time due to user behavior and network conditions. Ensuring that synthetic data models these temporal dynamics, including short-term fluctuations and long-term seasonality, is a major challenge.
High-dimensional Feature Spaces: Network traffic data is often high-dimensional, encompassing a wide range of features such as packet sizes, inter-arrival times, protocol flags, and source-destination IP pairs. Capturing correlations and dependencies across these features in a synthetic dataset is difficult. Synthetic data may either over-simplify these relationships, lose important interactions, or become too complex to be useful for downstream tasks.
Future Direction: Future commercial traffic generators could benefit from advanced generative techniques like DMs and Transformer-based architectures (GPTs), which, pre-trained on large datasets, offer powerful sequence generation capabilities. By bridging the gap between the LLM’s word embedding space and the embedding space of time-series tokens or packet flows, GPTs can be fine-tuned to generate realistic and contextually rich network traffic patterns. Similarly, DMs, known for their iterative refinement of data from noise, offer an exceptional ability to model intricate data distributions and generate synthetic data with high fidelity. These models could be effective in mimicking the statistical and temporal properties of network traffic, capturing both short-term variations and long-term dependencies.
Another way to capture complex patterns is through multimodal deep generative models that multiple data types (e.g., combine text, images, network metadata (e.g., headers, flow statistics), and raw RF signals). DMs such as Stable Diffusion can be conditioned on network topology diagrams or textual descriptions of attack scenarios to generate structured protocol-compliant synthetic traffic. Similarly, transformer (e.g. Contrastive Language-Image Pretraining (CLIP) [144]) models can align textual representation of network behavior with images or other modalities, allowing for more context-aware traffic synthesis that is more interpretable and adaptive. GPTs can extend their sequence modeling capabilities beyond packet flows by incorporating visual and textual network telemetry, enabling cross-modal learning for real-time network traffic simulation. The synergy of these techniques in commercial applications could enable scalable and precise traffic generation solutions, addressing long-standing challenges such as data imbalance, privacy preservation, and real-time traffic simulation. By leveraging these SOTA multimodal generative models, the next generation of commercial traffic generators could offer enhanced capabilities, transforming how synthetic network traffic is created and utilized for cybersecurity, network optimization, and anomaly detection.
@@ VII-D Evaluating Synthetic Data Quality
Challenge: Another key challenge lies in developing standardized metrics for evaluating the quality of synthetic network traffic data. While there are several metrics such as statistical similarity and performance on downstream tasks, there is no one-size-fits-all solution. There are many papers [82, 80, 157, 156] attempting to provide high-quality synthetic data, however, there are challenges yet to be solved.
Lack of Benchmark Datasets: Unlike other fields such as computer vision or NLP, there are relatively few publicly available, standardized benchmark datasets for synthetic network traffic evaluation. This makes it difficult to compare the performance of different synthesis approaches, especially for privacy-preserving techniques.
Generalization vs. Overfitting: Striking the right balance between generalization and overfitting in synthetic data generation remains a major issue. Overfitting to the real traffic can lead to privacy risks, while excessive generalization may result in synthetic data that fails to capture the essential characteristics of the network traffic, reducing its utility.
Task-Specific Evaluation: The effectiveness of synthetic data can vary significantly depending on the specific task it is used for, such as traffic classification, anomaly detection, or intrusion prevention. This calls for task-specific evaluation metrics, but defining these metrics can be complex and may require different evaluation frameworks for each application.
Future Direction: A unified evaluation toolkit combining statistical, task-specific, and privacy-preserving metrics is needed. Synthetic traffic benchmarks, similar to ImageNet [48] in the computer vision domain, would facilitate standardized and consistent comparisons across traffic generation methods. Also, interpretable metrics linking traffic properties with task-level performance can help building better generation models.
@@ VII-E Computational Complexity and Scalability
Challenge: As networks expand in size and complexity, the computational demands of generating and evaluating synthetic data grow accordingly. This also poses a particular challenge for real-time network traffic generation on edge devices, where computational resources are limited. To address this, the authors of [97] introduced lightweight DMs designed for efficient data generation on edge devices. However, further challenges remain in balancing computational efficiency with the accuracy and fidelity of the generated traffic.
@@
Scalability of Data Generation: Synthesizing data mimicking large scale networks (with sheer volume of data) requires significant computational resource. Many existing synthesis methods struggle to scale effectively, leading to bottlenecks in generating synthetic data for large network environments.
@@
Real-time Synthesis: In some applications, synthetic data may need to be generated in real-time to support testing and analysis under current network conditions. Achieving real-time synthesis is particularly challenging when attempting to replicate complex traffic patterns while maintaining high fidelity and privacy guarantees.
Future Direction: Future directions include the development of highly parallelized generation pipelines, federated learning and model distillation techniques for real-time inference. Edge-optimized models can be tailored to resource constraints while preserving core behavioral patterns. Integration into systems like Ericsson’s 5G DT [127], simulated in NVIDIA Omniverse, opens possibilities for real-time simulation and testing under realistic and evolving network conditions, supporting next-generation AI-driven network operations. Further, LLMs in many applications can also be replaced with evolving Small Language Models (SLMs).
VII-F Synthetic Data Generation for Unseen Network Scenarios
@@
Challenge: Synthesizing network data that represents unseen scenarios is crucial for evaluating system resilience and preparedness in various environments. Generative models can simulate unseen network conditions by learning from available datasets and extrapolating patterns for unobserved scenarios, such as extreme network loads, and rare cyberattacks etc. Additionally, when real-world data for specific platforms or protocols is unavailable, synthetic data generation can incorporate domain knowledge to approximate these environments. However, the key challenge lies in ensuring that the generated data remains both valid and useful for downstream tasks.
@@
Future Direction: One strategy is to follow Neural Style Transfer (NST) [60], which blends the content of one image with the style of another. NST is based on the intuition that the content (structure and high-level features) and style (textures, patterns, and colors) of an image can be extracted using different layers of a deep convolutional network and that optimizing a new image to match the content features of one image and the style features of another generates a result that retains the structure of the content image while adopting the artistic characteristics of the style image. To follow a similar approach, we first define two attributes for each network trace, Content which refers to the actual content/payload carried by the traffic trace (such as a video stream, or website), and Domain which refers to the conditions of the network which affects the traffic. For instance, when considering a network trace for a user streaming videos, the specific video title would be the content and the domain could be the specific network the user is connected to (wired or wireless) or the specific geographical location. The notation $T_{n}(C_{x},D_{y})$ represents a traffic trace corresponding to content $x$ and domain $y$. For example, consider a scenario where traffic traces with content $C_{1}$ from domain $D_{1}$, and traffic traces with content $C_{2}$ from domain $D_{2}$ are available, but traffic traces with content $C_{2}$ from domain $D_{1}$ is not available and needs to be generated. Analogous to NST, if we can extract content-specific features from $T_{2}(C_{2},D_{2})$ and domain-specific features from $T_{1}(C_{1},D_{1})$, these features can be combined to generate a new sample $T_{3}(C_{2},D_{1})$ representing $C_{2}$ and $D_{1}$.
While the layers of deep convolutional networks for extracting content and style in images are well established, no comparable approach has, to our knowledge, been studied for network traffic traces, leaving it an open research problem. However, P. Smolovic et al. [158] introduced the use of triplet loss to extract video-specific (content) features from network traffic traces, independent of the platform on which the video was streamed (domain). Based on this, we propose that using a methodology leveraging contrastive loss functions (e.g., triplet loss), to extract both content and domain-specific features from network traffic traces is feasible. Once the content and domain-specific features are extracted, the same approach as NST can be followed to generate the new sample by optimizing a random input to match the extracted content and domain.
An alternative is to use an encoder–decoder architecture such as a VAE, which structures the latent space into clusters by domain and content type. As shown in Figure 14, the model can form distinct domain clusters, with further subdivisions by content. Once trained, the latent vector for a traffic trace could be extrapolated in the latent space to generate a trace with a different content or domain. For example, Figure 14 shows a scenario where the model is trained on traces with content $C_{1}$ from both domains and $C_{2}$ from $D_{2}$ only. Accordingly, it will learn both $C_{1}$ and $C_{2}$ clusters in $D_{2}$ and only $C_{1}$ cluster in $D_{1}$. Once trained a trace $T(C_{1},D_{1})$ can be extrapolated in the latent space to create $C_{2}$ in $D_{1}$ to match context $C_{2}$ based on the relationship between $C_{1}$ and $C_{2}$ in $D_{2}$. Then the decoder can generate a realistic network traffic sample $T(C_{2},D_{1})$ that represents context $C_{2}$ and domain $D_{1}$.
VIII Conclusion
We presented a comprehensive review of synthetic network traffic generation using different mechanisms, mainly focusing on deep learning–based approaches. We discussed widely used statistical methods and commercial simulation tools for network traffic generation, as well as the different traffic types employed in the generation process. Furthermore, we provided a comparative analysis of these mechanisms through a novel approach based on LLMs. This structured classification offers insights into the strengths, limitations, and practical applications of each approach, showing the evolution of the field and how these methods are applied to synthesize diverse scenarios. Finally, we highlighted the existing challenges in network traffic generation and with future directions to guide the further research.
Several open challenges remain for the community, including balancing privacy with utility, adhering to protocol standards and realistic conditions, capturing diverse and complex behaviors, and developing standardized benchmarks. Scalability and efficiency for real-time synthesis, as well as generating realistic traffic for rare scenarios, are also pressing needs. Addressing these issues will drive the next wave of innovation, enabling more accurate, scalable, and trustworthy models for applications in cybersecurity, network optimization, and system testing.
IX Acknowledgement
This research was partially funded by the Australian Research Council Industrial Transformation Research Hub for Future Digital Manufacturing (IH230100013).